AnimeGhoul8863:

Peter played his favorite video game for 10 hours last week. Today, Peter's parents restricted him to 5 additional hours each week for the next 8 weeks. Create the function, f(x), that models Peter's total video game time, and explain what each number in the situation represents graphically, using complete sentences.

AnimeGhoul8863:

this one seems simple-er so imma try and see what i get

AnimeGhoul8863:

does this mean he started off with 10 hours a week to 15 a week cause it says 5 additional hours each week

AnimeGhoul8863:

@dude Help

AnimeGhoul8863:

im thinking x= Hours played while y=Weeks

AnimeGhoul8863:

@Vocaloid

AnimeGhoul8863:

still dont know am i adding or subtracting

Elsa213:

``` f(X) = 10 + 5X f(X) is the total hours for that week X is the week number after this plan was put in place Graphically: 10 is the y-intercept; this is when X=0 which is when the plan went into effect 5 is the change in the y-value for a change of 1 in the X-value; this will raise the number of games he can play by 5 each additional week ``` Source: https://www.wyzant.com/resources/answers/236959/explain_what_each_number_in_the_situation_represents_graphically_using_complete_sentences

dude:

Masta Elsa

Elsa213:

Pasta Dewd

AnimeGhoul8863:

so Y= Hours while X=Weeks ok so f(x)=10+5x so graphically it would look like this

Shadow:

Yep

Elsa213:

That Brainly tab, though. e.e

Shadow:

Feels betrayed

AnimeGhoul8863:

x'D i was trying to find the answer cause everyone went off line and that place sucks there is no answer

AnimeGhoul8863:

so i have that now how do i add the 8 weeks

Shadow:

You look at the graph and move over 8 units. Move up to where the line is at 8 units then note the y value.

Shadow:

From 0 -> 8 on the x axis

AnimeGhoul8863:

it lands on 50

Shadow:

Or just do math. Elsa gave you the function. X represents the weeks so substitute 8 for x then solve for y which is hours.

AnimeGhoul8863:

so f(x) = 10 + 5X x=8 y=50 hours in 8 weeks

AnimeGhoul8863:

50=10+5(8) ?

Shadow:

Yeah, you have checked your work

Shadow:

So now you need to explain what 10 and 5 are

AnimeGhoul8863:

10 is the original hours that was spent playing games while 5 is the new set of ours spent on playing games?

Shadow:

Graphically

AnimeGhoul8863:

oh umm let me think

AnimeGhoul8863:

10 equals the y intercept/hours played while 5 is the extra hours being played

AnimeGhoul8863:

?

Shadow:

You don’t explain 5 in terms of the graph correctly.

AnimeGhoul8863:

how

Shadow:

Just look at the graph, what is five?

AnimeGhoul8863:

O.O the slope

Shadow:

Yes

Shadow:

Lol those eyes

AnimeGhoul8863:

ok hold on let me write all dis down now

AnimeGhoul8863:

The function for Peters total game play is f(x)=10+5x. y= the amount of time being played while x=the weeks he has been playing. 10 equals the y intercept/hours being played while 5 equals the slope. So Graphically it would look like y=50 Hours in 8 weeks and x= 8 weeks.

AnimeGhoul8863:

^is it good

AnimeGhoul8863:

if so i have 1 last question and im done

AnimeGhoul8863:

@Shadow